OpenStudy (anonymous):

integral of 1/(x^2sqrt(4x+1))? I tried trig substitution since it looks like sqrt(x^2+a^2) with x=2sqrtx and a = 1

OpenStudy (anonymous):

x=a*tan (theta)

OpenStudy (anonymous):

I worked this down to 8*integral of [tan(theta)sec^2(theta)/[tan^4(theta)sec(theta)]

OpenStudy (anonymous):

simplify tangent and secant?

OpenStudy (anonymous):

is that 4x+1 inside, or 4x^2+1?

OpenStudy (anonymous):

I don't think trig sub works here like that, check your " it looks like sqrt(x^2-a^2) with x=2sqrtx and a = 1" idea again.

OpenStudy (anonymous):

4x+1, otherwise I think I could get it

OpenStudy (anonymous):

???, but (2sqrtx)^2 = 4x and 1^2 = 1.

OpenStudy (anonymous):

I guess its called radicals instead of trig sub in my math book, may have gotten terms confused.

OpenStudy (anonymous):

It was just this part: " sqrt(x^2-a^2)" with that minus sign there. yours has a +

OpenStudy (anonymous):

try use u =sqrt(4.x+1)

OpenStudy (anonymous):

sorry it should be a + sign

OpenStudy (anonymous):

raphael, then where is my du in the integral?

OpenStudy (anonymous):

partial fractions again

OpenStudy (anonymous):

no its wrong

OpenStudy (anonymous):

I was trying to get tan(theta) and sec^2(theta) in my integral so I could use u substitution, but I got sec/tan^3

OpenStudy (anonymous):